# Beyond the Silence: How Men Navigate Infertility Through Digital Communities and Data Sharing

### v1

@@ 3. Methods
3.1. Model Training and Parameter Tuning
We applied BERTopic (Grootendorst, 2022) to Reddit threads treated as documents, using BERT embeddings to cluster semantically similar content (Reimers and Gurevych, 2019; Angelov, 2020). To optimize performance we trained 32 models, selecting the best via u_mass coherence (Newman et al., 2010; Röder et al., 2015; Chang et al., 2009). The top configuration (coherence = –0.79) produced 115 topics. For each comment we assigned the dominant topic based on BERTopic’s probability distributions (McInnes et al., 2017; Campello et al., 2013). Full hyperparameter details appear in Appendix A.1.
@@ 3.1.1. Qualitative coding
3.2. Time-lagged Regression: Future Use Predictors
We analyzed comment-level records from r/maleinfertility, containing a user identifier and a timestamp. Events were binned into monthly periods. We operationalized community activity as the number of comments per user per month. To account for temporal dynamics, we created lagged predictors that reflected each feature’s value in the prior month. The dependent variable was defined as the number of comments contributed in the subsequent month ($t+1$). Only users with at least two months of data were included.
@@ 3.2.1. Feature Set Permutations
3.3. Studying interactions within the r/maleinfertility community:Network Analysis
We built a directed, weighted reply network where nodes are users and edges represent replies. Edges were aggregated by interaction count, self-loops removed, resulting in 8,536 nodes and 53,081 edges.
@@ 3.3.3. Detecting Echo Chambers
Following Morini et al. (Morini et al., 2021), we define echo chambers as clusters with Conductance ¡ 0.5 and Purity ¿ 0.7. To define purity, we start by labeling user stance which we defined as (Qualitative Theme — VADER sentiment). In other words, for each theme, a stance would be (theme X positive; theme X negative; and theme X neutral). This is in keeping with earlier work on stance as a product of topic and sentiment (ALDayel and Magdy, 2021; Naskar et al., 2016).
@@ 3.3.4. Qualitative Analysis of Clusters
For each Louvain cluster, we sampled 10 comments from the most central users (by PageRank), double-coded them, and met to reach consensus on cluster character and central user roles. Initial categories (mentor, moderator, expert) were refined to include nuanced behaviors exhibited by those roles such as system critic and newcomer.
3.4. Classifier: Who are the cross-posters?
@@
We used the same feature set described in detail in §3.2.1. We framed cross-posting (posting to both r/maleinfertility and r/infertility) as a binary classification task. Using the same 95 features, we trained and evaluated Logistic Regression, Random Forest, Extra Trees, Gradient Boosting, HistGradientBoosting, AdaBoost, GaussianNB, KNN, LinearSVC, and XGBoost/LightGBM with 5-fold StratifiedKFold. The best model achieved accuracy = 0.846, precision = 0.793, recall = 0.307, F1 = 0.442, and ROC–AUC = 0.835, indicating strong fit (Lemeshow and Hosmer Jr, 1982; Fawcett, 2006; Swets, 2014). We applied SHAP to the final Gradient Boosting model to identify the most predictive features (Lundberg and Lee, 2017). Each SHAP value indicates a feature’s additive effect on the log-odds of cross-posting, with positive values increasing and negative values decreasing the likelihood of cross-posting (Lundberg and Lee, 2017).
@@ 3.5. Positionality and Ethical Stance
We are a team of researchers committed to supporting individuals navigating trauma, uncertainty, and life transitions. The first author, a cisgender man, studies how social technologies empower vulnerable communities while posing risks to privacy, autonomy, and identity. The second author, a cisgender South Asian woman, is a graduate student focused on reproductive and public health. The third author, a cisgender East Asian woman, examines how digital technologies can empower users. The last author, a nonparent white Latinx cisgender woman, specializes in transforming large, complex datasets into actionable insights for nonprofits, government, and industry, emphasizing the human stories behind the data.
Aware that online users often do not expect to be quoted directly (Fiesler and Proferes, 2018; Fiesler et al., 2024), we paraphrased community stances and edited supporting quotes to protect anonymity (Bruckman, 2002).
@@ 4. Findings
Our findings unfold across three layers of analysis. In §4.1, we present a thematic analysis showing how medical advice, emotional support, and moderation position the subreddit as a diagnostic hub, therapeutic commons, and digital institution, and how these practices influence future engagement. In §4.2, we examine community structure and roles along with how moderation and partner participation shape collective identity. Finally, §4.3 explores cross-posting, revealing how users who bridge r/maleinfertility and r/infertility serve as navigators and mentors, aligning norms across spaces and reinforcing male-centered support.
4.1. Diagnostic Utility Over Emotional Support: Themes Driving Future Engagement
Our analysis of 115 discussion topics, grouped under three overarching themes—Medical Advice, Emotional Support, and Moderation—illustrates how the infertility community on Reddit blends biomedical knowledge, affective support, and digital governance to sustain its members. Each theme encompasses multiple sub-themes that together reflect the community’s hybrid role as an informal medical forum, a therapeutic space, and a structured support system. Medical Advice (63.3%) centers on interpreting diagnostic tests, navigating treatment pathways, confronting advanced biological challenges, and addressing systemic barriers such as insurance. Emotional Support (7.4%) encompasses grief, coping rituals, comfort posting, and strategies for fostering resilience and social connection. Moderation (29.2%) highlights on-boarding rituals, navigation aids, rule enforcement, and the wiki/FAQ infrastructure that structures participation. Together, these themes show how the community functions simultaneously as an informal diagnostic hub, a therapeutic commons, and a regulated digital institution. A more detailed breakdown of the themes and topics can be seen in Appendix §A.2.
@@ 4.1.1. Medical Advice: Navigating Diagnostics, Treatments, and Systemic Barriers
The most prominent theme centers on medical advice and information exchange, where users collectively interpret results, evaluate treatment options, and contextualize both biological and systemic challenges.
@@
A large cluster of discussions revolves around decoding semen analyses and other fertility test reports. Posters often admit confusion about terms like motility, morphology, or DNA fragmentation, with some worrying that a flagged number meant they were “doomed.” Peers step in to explain contested measures, reframing results as part of a broader diagnostic picture. For example, users emphasized that “percentages can mislead if counts are high,” reminding others that even a 0% morphology result did not rule out conception.
Female Infertility Generalized Guidance
Parallel threads focus on ovulation induction, endometrial lining preparation, or embryo transfer logistics. Posters compare regimens of Clomid, Letrozole, or estrace, and narrate experiences of thin lining or cycle cancellation. Discussions often reveal systemic frustrations: one user described “feeling like fertility treatment is another full-time job” due to scheduling battles and unclear protocols.
Male Infertility and Biological Limits
Discussions of azoospermia, low motility, and donor sperm highlight how male infertility intersects with identity and grief. Some describe the devastation of failed surgical retrievals, while others detail the emotional weight of using donor gametes—one poster admitted being more anxious about “telling family than actually using donor sperm.”
@@
Beyond treatment, members share lifestyle strategies: supplements such as CoQ10, vitamin E, and L-carnitine, alongside advice to “ditch the laptop on the lap” or “switch to cooler underwear.” While highly anecdotal, these exchanges position men and women as active agents in shaping outcomes.
@@
Another critical sub-theme is financial navigation. Posters swap tips on discount programs (e.g., Compassionate Care), employment benefits (e.g., Starbucks part-time IVF coverage), and shared-risk refund schemes. Others compare international systems, noting for example that “in the Netherlands, three cycles are covered with a deductible” whereas U.S. cycles often cost over \$10,000 out-of-pocket. This highlights inequities where access to Assistive Reproductive Technology (ART) depends as much on geography and insurance literacy as biology. When semen is only viable for analysis within several hours of taking a sample, users debated the merits of various brands of at-home testing kits when access to clinics (along with the discomfort of providing in-person samples) was limited.
Taken together, these sub-themes reveal Reddit as a secondary diagnostic and logistical space: members crowdsource interpretations, trade experiential knowledge, and provide rapid responses that often fill gaps left by inaccessible or opaque medical care.
@@ 4.1.2. Emotional Support: Coping, Resilience, and Connection
Running parallel to biomedical exchange is a robust theme of emotional support, where the subreddit functions as a therapeutic scaffold.
@@
Threads often feature raw narratives following miscarriage, failed embryo transfers, or prolonged infertility. Posters describe feeling “numb,” “burned out,” or like they were “going through the motions of injections while emotionally checked out.” Community replies emphasize validation—reminding one another that every loss, even a chemical pregnancy, “is real and deserves mourning.” It is important to note that men who decide to have children with their partners are also expectant parents, sharing the full spectrum of emotions—joys and sorrows—throughout the process just as much as women do. These comments highlight that men feel the pain of infertility and the loss of a biological child just as deeply as their female counterparts.
@@
Alongside grief are lighter posts that serve as comfort rituals. Members trade injection playlists, joke about “pin the tail on the donkey with my thighs,” or celebrate small victories like mastering self-administered shots. These threads blend humor with encouragement, reframing difficult procedures as milestones of resilience. However, some posts suggested that the poster should share their challenges with their therapist in addition to a show of some support.
@@
The community also carves out space for tentative joy. Ritualized language such as “cautious congrats” during early betas tempers hope with shared recognition of risk. Others celebrate birthdays or procedural wins with “you guys rock!” affirmations, offering reprieve from clinical stress.
Talking to Family and Friends
A recurring pain point is how to navigate conversations with relatives and peers. Posters vent about insensitive pregnancy announcements or family pressure, while peers strategize about boundary-setting and selective disclosure. Advice often centers on protecting one’s “sanity first,” even if it means declining baby showers.
@@ 4.1.3. Moderation: Structure, Orientation, and Safety
Automated bots and moderators enforce strict posting guidelines, regularly reminding users that “standalone culture is saved for complex topics.” Basic questions, introductions, or updates are redirected into daily or weekly threads. Some users express gratitude for this structure, noting that the rules keep the subreddit “focused and high-signal,” avoiding the clutter of repetitive or low-effort posts. Others, however, describe the system as alienating when they first join. One poster remarked that their message was “not visible until a mod approved it,” leaving them feeling like they were “posting into a void” at a moment of emotional urgency.
@@
Navigation challenges are a recurrent frustration. Newcomers often post in the wrong place and are redirected with templated bot messages pointing to the FAQ, wiki, or Welcome Wednesday thread. For some, these messages serve as useful orientation tools that provide quick access to resources; others describe them as “robotic and impersonal,” especially when received after disclosing sensitive struggles. A common critique was that automated responses “felt more like bureaucracy than support,” creating disorientation during moments of crisis. Yet, once acclimated, many members acknowledge that the structured system of daily treatment chats and thematic threads helps them find relevant discussions and reduces the chaos of a large community. This navigation is not only done using bot, but at times members of the community act as navigators as well.
Onboarding and Private Messaging
Onboarding is a blend of humor, empathy, and procedural instruction. Automated greetings often reference the subreddit as “the best shitty corner of the internet,” a self-aware attempt to soften the rigidity of rule enforcement. Many users appreciate this levity, interpreting it as a sign that moderators understand the emotional intensity of infertility. At the same time, others note that such humor does not offset the sting of rejection when their posts are removed. In some cases, moderators also encourage members to continue sensitive or off-topic conversations via private messages, a practice that both protects thread coherence and provides space for more personal exchanges. While some valued this redirection as an invitation to connect more deeply, others felt it reinforced exclusion by pushing vulnerable disclosures out of public view.

### v2

@@ 3. Methods
3.1. Topic Model Training and Parameter Tuning (RQ1a)
To answer RQ1a, We applied BERTopic (Grootendorst, 2022) to Reddit threads treated as documents, using BERT embeddings to cluster semantically similar content (Reimers and Gurevych, 2019; Angelov, 2020). To optimize performance we trained 32 models, selecting the best via u_mass coherence (Newman et al., 2010; Röder et al., 2015; Chang et al., 2009). The top configuration (coherence = –0.79) produced 115 topics. For each comment we assigned the dominant topic based on BERTopic’s probability distributions (McInnes et al., 2017; Campello et al., 2013). Full hyperparameter details appear in Appendix A.1.
@@ 3.1.1. Qualitative coding
3.2. Time-lagged Regression: Future Use Predictors (RQ1b)
To answer RQ1b, we analyzed comment-level records from r/maleinfertility, containing a user identifier and a timestamp. Events were binned into monthly periods. We operationalized community activity as the number of comments per user per month. To account for temporal dynamics, we created lagged predictors that reflected each feature’s value in the prior month. The dependent variable was defined as the number of comments contributed in the subsequent month ($t+1$). Only users with at least two months of data were included.
@@ 3.2.1. Feature Set Permutations
3.3. Studying interactions within the r/maleinfertility community:Network Analysis (RQ2)
To answer RQ2, we built a directed, weighted reply network where nodes are users and edges represent replies. Edges were aggregated by interaction count, self-loops removed, resulting in 8,536 nodes and 53,081 edges.
@@ 3.3.3. Detecting Echo Chambers
Following Morini et al. (Morini et al., 2021), we define echo chambers as clusters with Conductance ¡ 0.5 and Purity ¿ 0.7. To define purity, we start by labeling user stance which we defined as (Qualitative Theme — VADER sentiment). In other words, for each theme, a stance would be (theme X positive; theme X negative; and theme X neutral). This is in keeping with earlier work on stance as a product of topic and sentiment (ALDayel and Magdy, 2021; Naskar et al., 2016). We used this method to answer RQ2a.
@@ 3.3.4. Qualitative Analysis of Clusters
For each Louvain cluster, we sampled 10 comments from the most central users (by PageRank), double-coded them, and met to reach consensus on cluster character and central user roles. Initial categories (mentor, moderator, expert) were refined to reflect nuanced behaviors exhibited by those roles such as system critic and newcomer. We used this method to answer RQ2b.
3.4. Classifier: Who are the cross-posters? (RQ3)
@@
We used the same feature set described in §3.2.1. We framed cross-posting (posting to both r/maleinfertility and r/infertility) as a binary classification task. Using the same 95 features, we trained and evaluated Logistic Regression, Random Forest, Extra Trees, Gradient Boosting, HistGradientBoosting, AdaBoost, GaussianNB, KNN, LinearSVC, and XGBoost/LightGBM with 5-fold StratifiedKFold. The best model achieved accuracy = 0.846, precision = 0.793, recall = 0.307, F1 = 0.442, and ROC–AUC = 0.835, indicating strong fit (Lemeshow and Hosmer Jr, 1982; Fawcett, 2006; Swets, 2014). We applied SHAP to the final Gradient Boosting model to identify the most predictive features (Lundberg and Lee, 2017). Each SHAP value indicates a feature’s additive effect on the log-odds of cross-posting, with positive values increasing and negative values decreasing the likelihood of cross-posting (Lundberg and Lee, 2017). This method allowed us to answer RQ3.
@@ 3.5. Positionality and Ethical Stance
Our research team is committed to supporting individuals navigating trauma, uncertainty, and life transitions. The first author, a cisgender man, occupies partial insider status—sharing gender identity with participants while maintaining analytical distance—and studies how social technologies simultaneously empower vulnerable communities and pose risks to privacy, autonomy, and identity. The second author, a cisgender South Asian woman and graduate student in reproductive and public health, grounds the study in clinical and epidemiological context essential for interpreting community discussions of diagnostics and treatment. The third and fourth authors—cisgender East Asian and white Latinx women respectively—bring critical external perspectives on masculinity construction; the fourth author’s specialization in surfacing human stories behind complex datasets bridges our computational and qualitative methods.
Following AoIR Ethical Guidelines (Franzke et al., 2020), we mitigated risks of weaponization by bad actors. Because direct quotes might contain sensitive information usable against individuals (Sloan and Quan-Haase, 2022), and online users often do not expect to be quoted in research (Fiesler and Proferes, 2018; Fiesler et al., 2024), we paraphrased community stances and edited supporting quotes to protect anonymity (Bruckman, 2002; Markham, 2012), preventing digital traceability while preserving original intent. Disguised quotes appear italicized throughout the paper for clarity.
@@ 4. Findings
Our findings unfold across three layers of analyses. In §4.1, we answer RQ1a by presenting a thematic analysis showing how medical advice, emotional support, and moderation position the subreddit as a diagnostic hub, therapeutic commons, and digital institution. We go on to answer RQ1b in §4.2 by describing what influenced future community engagement. In §4.3, to answer RQ2, we examine community structure as well as user roles—mentors, experts, hope brokers, and system critics—along with how moderation and partner participation shape collective identity. Finally, to answer RQ3, §4.5 explores cross-posting, revealing how users who bridge r/maleinfertility and r/infertility serve as navigators and mentors, aligning norms across spaces and reinforcing male-centered support.
4.1. Diagnostic Utility Over Emotional Support: Themes Driving Future Engagement (RQ1a)
Our analysis of 115 discussion topics, grouped under three overarching themes—Medical Advice, Emotional Support, and Moderation—illustrates how the infertility community on Reddit blends biomedical knowledge, affective support, and digital governance. Medical Advice (63.3%) centers on interpreting diagnostics, navigating treatment, confronting biological challenges, and addressing systemic barriers like insurance. Emotional Support (7.4%) encompasses grief, coping rituals, comfort posting, and strategies for resilience. Moderation (29.2%) highlights onboarding, navigation aids, rule enforcement, and wiki/FAQ infrastructure. Together, these themes reveal how the community functions simultaneously as an informal diagnostic hub, therapeutic commons, and regulated digital institution—providing an antidote to the isolation many men experience during infertility. A detailed breakdown appears in Appendix §A.2.
@@ 4.1.1. Medical Advice: Navigating Diagnostics, Treatments, and Systemic Barriers
The most prominent theme centers on medical advice and information exchange, where users interpret results, evaluate treatments, and contextualize biological and systemic challenges. Given male infertility’s diverse causes, discussions focus on problem-solving: how do I overcome this specific problem?
@@
Facing research gaps that systematically exclude men from fertility studies (Harlow et al., 2020), users described ”grasping at straws” for male-factor information. Peers filled this void by collaboratively decoding semen analyses (Gannon et al., 2004), reassuring those who feared flagged numbers meant they were ”doomed” by reframing metrics contextually—explaining that percentages mislead when counts are high and that 0% morphology (shape and size of semen) [^1] Sperm morphology refers to the size and shape of sperm, which is one factor analyzed in a semen test to evaluate male fertility – Source: https://www.mayoclinic.org/diseases-conditions/male-infertility/expert-answers/sperm-morphology/faq-20057760 does not preclude conception.
@@
Beyond treatment, members share lifestyle strategies: supplements like vitamin E as well as advice to “ditch the laptop on the lap” or “switch to cooler underwear.” Though anecdotal, these exchanges position participants as active agents. Some share specific routines—workout timing, masturbation frequency, diets—crowdsourcing and normalizing conception efforts.
@@
Financial navigation forms another critical sub-theme. Posters swap tips on discount programs (e.g., Compassionate Care), employment benefits (e.g., Starbucks part-time IVF coverage), and shared-risk refund schemes. Others compare international systems: “in the Netherlands, three cycles are covered with a deductible” versus U.S. cycles often exceeding \$10,000 out-of-pocket. This highlights inequities where access depends on geography and insurance literacy as much as biology. Users advocated for at-home testing kits when clinic access was limited, but this self-advocacy was complicated by labs repeatedly rejecting samples. One user lamented the discouraging cycle of awkwardly producing multiple specimens only to face continued rejection.
(Fe)male Infertility Generalized Guidance: Navigating Infertility and Partner Dynamics
These sub-themes reveal Reddit as a secondary diagnostic and logistical space where members crowd-source interpretations and trade experiential knowledge, filling gaps left by inaccessible medical care. Threads span female concerns—ovulation, endometrial preparation, embryo transfer—and male diagnoses like azoospermia [^2] Condition where a man’s ejaculate contains no sperm — Source https://my.clevelandclinic.org/health/diseases/15441-azoospermia and low motility. One user described treatment as “another full-time job”; others expressed frustration with medications improving hormone levels while sperm counts remained zero—“heartbreaking to fix the chemistry on a chart but still see no change.” Yet these informational exchanges carry emotional weight. Partners describe male factor infertility transforming “my problem into our battle,” with one writing: “I would rather have this husband and no baby than a baby and no husband.” Men echoed this, framing relationships as a “trust fall”—partners must support each other’s grief without hiding their own pain. Some tensions emerged, with moderators asking partners to avoid being “all doom and gloom” and stick to medical facts, recognizing men already face significant pressure. As these examples suggest, medical discussions often blend into emotional territory—a theme we explore more fully in the following section.
@@ 4.1.2. Emotional Support: Coping, Resilience, and Connection
Running parallel to biomedical exchange is a robust theme of emotional support, where the subreddit functions as a therapeutic scaffold. Members here actively provide emotional care work (see §2.2.2) such as acknowledging that dealing with ”miscarriages and infertility is miserable - all of your feelings are valid.” Words of encouragement to ”accept yourself in your disease” but not be ”defined” by it are often repeated. Newcomers are encouraged to ask many questions as they want, and to DM and reach out to veteran members of the community.
@@
Threads feature raw narratives following miscarriage, failed transfers, or prolonged infertility. Posters describe feeling ”numb,” ”burned out,” or ”going through the motions while emotionally checked out.” Community replies emphasize validation—reminding members that every loss, even a chemical pregnancy, [^3] Pregnancies ending early, around the fifth week – Source: https://my.clevelandclinic.org/health/diseases/22188-chemical-pregnancy ”is real and deserves mourning.” Central to this grief is confronting that one may never have a biological child—whether through failed surgical retrievals [^4] Surgical sperm retrieval procedure that uses an operating microscope to collect mature sperm for use in in vitro fertilization (IVF) - Source: https://tinyurl.com/y3uranhz or the emotional weight of donor gametes. [^5] Gametes are the reproductive cells (sperm in males and eggs/ova in females) – Source: https://code-medical-ethics.ama-assn.org/ethics-opinions/gamete-donation Yet alongside grief, members affirm that ”our family is of our own making,” declaring themselves fathers regardless of biological link:
Deep down, guys often equate masculinity with fertility, so using a donor can feel like it defies our basic programming. That’s why partners must validate that we aren’t ‘lesser men’ because of a medical issue. I’ve learned that fatherhood isn’t about DNA or seeing my features in my child’s face—it’s seeing my influence in her character. Guiding her through milestones and modeling how she should be treated is what truly makes her my daughter.
Talking to Family and Friends
A recurring pain point is navigating family conversations. Posters describe gatherings as emotional minefields—fielding intrusive questions, insensitive announcements, or comments like ”it’s different when you fathered the child.” Those using donor conception describe imposter syndrome as fathers; one admitted ”more anxiety about telling family than actually using donor sperm,” while another wrote that accepting congratulations ”felt like I was misleading the family” and baby showers were ”laden with anxiety because I knew it wasn’t me who got my wife pregnant.” The pain intensifies with tragedy—one poster described repeatedly explaining to thrilled family, including an ailing parent, that ”there were no heartbeats and the babies were gone,” turning their support system into ”a source of exhaustion.” Peers strategize about boundary-setting and selective disclosure, advising protection of one’s ”sanity first” even if it means declining events.
@@
Alongside grief are lighter posts serving as comfort rituals. Members deploy pop culture references to celebrate good results, joke about alternative treatments and past choices now laden with irony, and describe teasing partners ”mercilessly and lovingly” to weather slumps together. These threads blend humor with encouragement, reframing difficult procedures—from mastering self-administered testosterone injections to facing surgical sperm extractions—as milestones of resilience. Beyond peer support, members actively recommend professional help; one noted that couples therapy ”helped a ton” and emphasized one need not reach a breaking point to benefit, while another urged a struggling user to see a counselor for better coping strategies.
@@
The community carves out space for tentative joy while navigating complex emotional boundaries. Ritualized language like ”cautious congrats” tempers hope with shared recognition of risk, and success stories often appear behind spoiler warnings to protect those not ready to see them. Members celebrate vicariously—one user described another’s post-surgery baby as ”the best thing I could’ve read today,” finding hope that ”there is still a chance.” and affirmations like ”you guys rock!” offer reprieve from clinical stressors. Yet users must navigate implicit ”pain olympics” rules governing whose suffering warrants voice; some newcomers find this emotional intensity intimidating, encountering harsh policing rather than support—a tension we explore by analyzing moderation next.
@@ 4.1.3. Moderation: Structure, Orientation, and Safety
Onboarding and Private Messaging
Onboarding is a blend of humor, empathy, and procedural instruction. Automated greetings often reference the subreddit as “the best shitty corner of the internet,” a self-aware attempt to soften the rigidity of rule enforcement. Many new users appreciate this levity, interpreting it as a sign that moderators understand the emotional intensity of infertility. Often newcomers are greeted with an emphatic ”Welcome!” as well as an invitation to refer to community pinned resources but to not shy away from asking questions. Despite that, some note that such humor does not fully offset the sting of rejection when their posts are removed for not adhering to group rules. In some cases, moderators also encourage members to continue sensitive or off-topic conversations via private messages, a practice that both protects thread coherence and provides space for more personal exchanges. While some valued this redirection as an invitation to connect more deeply, others felt it reinforced exclusion by pushing vulnerable disclosures out of public view. Here again we can see signs of TMI as exemplified by the limited autonomy that new members can express themselves in this subreddit (Englar-Carlson, 2006; Agüera et al., 2017).
@@
Often part in parcel of the newcomer experience might be meeting with their posts automodded and removed. Automated bots and moderators enforce strict posting guidelines, regularly reminding users that they enforce keeping to the topic of male infertility; ”reading the room” is one phrase used to make new posters aware that outside discussions are not welcome. Even basic questions, introductions, or updates are redirected into daily or weekly threads. Automated mods are not the only enforcers of these policies. Seasoned participants remind posters that this subreddit serves men with clinical infertility—not adjacent concerns like low libido, general hormone optimization, [^6] Maintaining hormones within optimum ranges becomes more difficult with age. Though medically contested (Holtorf, 2009), it is discussed in medical resources: https://nyulangone.org/news/future-you-how-hormones-impact-health-every-age or reversing steroid use. Members also distinguish clinical infertility from general trying to conceive (TTC): users with normal parameters, or those trying for under a year without severe issues, are redirected to TTC subreddits as their situation lacks the medical and emotional gravity of diagnosed infertility. One moderator drew sharp distinctions: between ”actual issues” and self-inflicted consequences of steroid use warranting reality checks rather than support; between the ”agony,” ”guilt and shame” of men facing infertility versus requests for hormone optimization ”coaching”; and between struggling to conceive versus ”odd threads” about contraception or sexual mechanics. These off-topic discussions, members emphasized, can be ”triggering” for participants navigating genuine loss—hence their strong discouragement, often with expletives.
Many users express gratitude for this structure, noting that the rules keep the subreddit “focused and high-signal,” avoiding the clutter of repetitive or low-effort posts. Others, however, describe the system as alienating when they first join. One poster remarked that their message was “not visible until a mod approved it,” leaving them feeling like they were “posting into a void” at a moment of emotional urgency. Though the very fact that the rules can be discussed within the forum demonstrate that - to a certain degree - they are flexible and adaptable. In our research, we did encounter users providing feedback on what structures of posts made more sense if ”there wasn’t enough traffic for recurring posts” because they ”are too busy on other parts of Reddit.” Managing the flow of information does seem to be a contested topic open for debate.
@@
As is evident from the above themes, navigating where the flow of conversations are structure to happen can be a recurrent frustration. Newcomers often post in the wrong place and are redirected with templated bot messages pointing to the FAQ, wiki, or Welcome Wednesday thread. For some, these messages serve as useful orientation tools that provide quick access to resources; others describe them as “robotic and impersonal,’’ especially when received after disclosing sensitive struggles. A common critique was that automated responses “felt more like bureaucracy than support,” creating disorientation during moments of crisis. Yet, once acclimated, many members acknowledge that the structured system of daily treatment chats and thematic threads helps them find relevant discussions and reduces the chaos of a large community. This navigation is not only done using bots, but at times members of the community act as navigators as well. By pointing newcomers to ”the FAQ at the top of the page” or ”pinned posts” that contains the experiential brain trust developed by members over the years, they try to assist in making sure ”most of your questions” get answered in a timely manner.
r/infertility vs. r/maleinfertility
Members articulated several perceived differences between r/infertility and
r/maleinfertility, though these characterizations were contested. First, demographics: r/infertility was described as female-dominated, where men are ”rare” and the conversation feels female-focused, while r/maleinfertility offers space for men to discuss how societal emphasis on virility shapes their experience without the ”ban on men showing weakness.” Second, focus: some perceived r/infertility as primarily emotional, with one user finding it ”intimidating” and another reporting judgmental responses to scientific questions about lifestyle factors. By contrast, r/maleinfertility was characterized as practical and data-driven—members engage in technical discussions of semen analysis parameters and treatment protocols. It is important to note that these distinctions were not universally held. One female member of r/infertility pushed back, clarifying that members there are ”very much interested in the scientific aspects”—suggesting the emotional-versus-scientific framing oversimplifies both communities.
4.2. Predictors of More Future Engagement (RQ1b)
The time-lagged regression analysis presented in Tables 1-3 offers insights into how language use (sentiment), themes discussed by the user, and user engagement behavior on this subreddit community predict higher engagement from month to month in the infertility support community. Specifically, linguistic markers drawn from LIWC (Table 1) show that future engagement and language showing affiliation with others in the community, insight-oriented words, and cognitive processing terms are all positively associated with an increased engagement in the community. Conversely, language associated with detachment or externalization—such as references to money, negation, or fatalistic expressions—is negatively associated with continued participation.